OpenStudy (anonymous):

#riddle 5 What is half of 2+2?

OpenStudy (hennawho):

3

OpenStudy (anonymous):

3

OpenStudy (darthvader2900):

3

OpenStudy (scienceteen8):

3

OpenStudy (darthvader2900):

lol

OpenStudy (alexandervonhumboldt2):

wrong: (2+2)/2=2

OpenStudy (alexandervonhumboldt2):

\[\frac{ 2+2 }{ 2 }=2\]

OpenStudy (anonymous):

no they are right 3 (half of 2 is 1, and then 1 plus 2 is 3)

OpenStudy (andrewishere):

That's good lol. It's either asking 1/2 of 2, then add 2, which would equal 3 OR its is asking 1/2 of (2+2), which would be 2.

OpenStudy (andrewishere):

It is a riddle so I would say 3

OpenStudy (andrewishere):

A truck question

OpenStudy (andrewishere):

Trick

OpenStudy (alexandervonhumboldt2):

the question says: what is half of 2+2, not what is half of 2 plus 2

OpenStudy (alexandervonhumboldt2):

2

OpenStudy (andrewishere):

But you might be reading it wrong. It might be asking half of 2, then +2

OpenStudy (anonymous):

@AlexandervonHumboldt2 how was my post spam

OpenStudy (alexandervonhumboldt2):

because it is not related to math http://openstudy.com/code-of-conduct

OpenStudy (alexandervonhumboldt2):

If you feel that you warning was undeserved you can waste your time to contact abuse@openstudy.com

OpenStudy (anonymous):

well im sorry i am new to this my sis showed me open study...... u are not very helpfull

OpenStudy (andrewishere):

Wait so this was fake? That's not cool